OpenStudy (anonymous):

Let f(x) = x2 - 16. Find f-1(x) do i have to do the inverse on this to?

OpenStudy (anonymous):

Same idea as before, your function is y = x^2 - 16 Inverse Function is x = y^2 - 16 Lets solve for y: add 16 to both sides, you get x + 16 = y^2 So what is y = ?

OpenStudy (anonymous):

well?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

16